OpenStudy (anonymous):

prove that x^4-y^4=b^2-a^2, if a cot(theta)+b cosec(theta)= x^2 and b cot(theta) + a cose(theta)= y^2 ...??????

OpenStudy (anonymous):

sm1 help

OpenStudy (akashdeepdeb):

Tell me if you understand this or not. :)

OpenStudy (psymon):

\[x ^{4}=(acot \theta +b \csc \theta)^{2}\] \[y ^{4}=(acsc \theta + b \cot \theta)^{2}\] \[x ^{4}=a^{2}\cot^{2} \theta + abcot \theta \csc \theta + b^{2}\csc^{2} \theta \] \[y^{4} = a^{2}\csc^{2} \theta + abcot \theta \csc \theta + b^{2}\cot \theta\] x^4 - y^4 = \[a^{2}\cot^{2} \theta + abcot \theta \csc \theta + b^{2} \csc^{2} \theta - a^{2}\csc^{2} \theta - abcot \theta \csc \theta - b^{2}\cot^{2} \theta\] becomes: \[a^{2}\cot^{2} \theta - a^{2} \csc^{2} \theta+b^{2}\csc^{2} \theta - b^{2}\cot^{2} \theta\] \[a^{2}(\cot^{2} \theta - \csc^{2} \theta ) + b^{2}(\csc^{2} \theta - \cot^{2} \theta)\] Pythagorean identity says 1 + cot^2 = csc^2 Therefore cot^2(theta) - csc^2(theta) = -1 and csc^2(theta) - cot^2(theta) = 1 \[a^{2}(-1) + b^{2}(1) \implies b^{2} - a^{2}\] Therefore \[x^{4} -y^{4} = b^{2} - a^{2}\]

OpenStudy (psymon):

And I have no way to do work on paper and take a pic like akash, lol. Taking reliable pictures of anything is not something I can do x_x

OpenStudy (anonymous):

:D.... thnx anyway @Psymon .......and thnx @AkashdeepDeb ....i understud

OpenStudy (akashdeepdeb):

:) He DESERVES the medal. :)

OpenStudy (psymon):

Lol, you don't need my gibberish then.

OpenStudy (psymon):

Why do I?

OpenStudy (akashdeepdeb):

LaTeX is tougher than writing. :P

OpenStudy (psymon):

I don't have the ability to write, I have to use latex xD

OpenStudy (anonymous):

i knw......

OpenStudy (anonymous):

wish i could give bth of u a medal...... :D

OpenStudy (psymon):

He answered it just as well if not better and faster, I don't deserve any medal, lol.